OpenStudy (anonymous):

Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5).

OpenStudy (anonymous):

f(x) = 3x - 1 and g(x)= -x + 6, i.e. f(-2) +g(5) = 3(-2) - 1+ -5 + 6=-6-1-5+6=-6 i.e. f(-2) +g(5) = -6 is the required answer. @Jmat1212

OpenStudy (anonymous):

aww i see what i did wrong the last step